### v2

# SplashNet: Split‑and‑Share Encoders for Accurate and Efficient Typing with Surface Electromyography

Abstract
Surface electromyography (sEMG) at the wrists could enable natural, keyboard‑free text entry, yet the state‑of‑the‑art emg2qwerty baseline still misrecognizes 51.8% of characters zero‑shot on unseen users and 7.0% after user‑specific fine‑tuning. We trace much of these errors to mismatched cross‑user signal statistics, fragile reliance on high‑order feature dependencies, and the absence of architectural inductive biases aligned with the bilateral nature of typing. To address these issues, we introduce three simple modifications: (i) Rolling Time Normalization which adaptively aligns input distributions across users; (ii) Aggressive Channel Masking, which encourages reliance on low‑order feature combinations more likely to generalize across users; and (iii) a Split‑and‑Share encoder that processes each hand independently with weight‑shared streams to reflect the bilateral symmetry of the neuromuscular system. Combined with a five‑fold reduction in spectral resolution ($3\rightarrow$6 frequency bands), these components yield a compact Split-and-Share model, SplashNet‑mini, which uses only ¼ the parameters and 0.6× the FLOPs of the baseline while reducing character error rate (CER) to 36.4% zero‑shot and 5.9% after fine‑tuning. An upscaled variant, SplashNet (½ parameters, 1.15× FLOPs of the baseline), further lowers error to 35.7% and 5.5%, representing 31% and 21% relative improvements in the zero-shot and finetuned settings, respectively. SplashNet therefore establishes a new state-of-the-art without requiring additional data.
1 Introduction
Translating neuromuscular signals into typed text is a novel but rapidly developing area at the intersection of human-computer interaction and machine learning. The emg2qwerty dataset (14) is the first large-scale benchmark specifically for wrist EMG-based touch typing. It contains over 346 hours of data from 108 users typing sentences while wearing electrode bands on each wrist. This dataset was motivated by the promise of wrist EMG as an always-available input modality for AR/VR glasses and other scenarios where traditional keyboards are impractical. In parallel, related benchmarks like emg2pose have been developed for EMG-based hand pose estimation (13), reflecting a broad interest in neuromotor interfaces for controlling virtual objects, robots, or text entry. Early explorations of EMG for text input date back at least two decades: for example, Yu et al. (22) demonstrated an “EMG keyboard” that allowed a forearm amputee to input characters via muscle signals. These efforts were limited in scale and accuracy, but they proved the principle that muscle activity alone can convey typing information. Today, with much larger datasets and deep learning models, the accuracy of EMG-based typing has greatly improved (2; 14; 13), making it a viable research direction for assistive technology and next-generation user interfaces.
Although promising, EMG interfaces face critical challenges in generalization. EMG interfaces suffer from substantial domain shift across users and sessions: anatomy, electrode placement, fatigue, and day-to-day physiology all alter the signal for the same action (14). The emg2qwerty benchmark probes this shift with (i) a zero-shot test, where one trains on many users and tests on a new one, and (ii) a personalization test that finetunes on a few target-user samples. Their baseline ASR-style system (CNN on spectrograms + LM decoding) misrecognized $51.8\%$ of characters zero-shot, but finetuning cut errors to approximately $7\%$, showing that most residual errors are systematic, user-specific variations. Yet even large training pools struggle: Sivakumar et al. (14) estimate $\mathcal{O}(10^{3})$ users are needed for robust out-of-the-box performance, a trend echoed by CTRL-Labs’ gains from massive data (2).
While scaling may significantly increase the performance of EMG interfaces, collecting these large-scale datasets is time-consuming and human-resource expensive. We instead present a simpler and complementary path towards improving sEMG generalization. Our key contribution is to first identify limitations in sEMG data that form obstacles to zero-shot generalization. We then propose simple, causal, and computationally inexpensive modifications to address these limitations, resulting in models that are more invariant to user/session differences. Our rationale for pursuing causal and computationally inexpensive modifications is to enable sEMG decoders to work on-device, an important consideration for future wrist EMG devices.
2 Insights into improving sEMG decoding
We summarily make three insights, and subsequently, three simple modifications to sEMG decoding that substantially improve zero-shot and finetuning performance while reducing computational costs.
Figure 1: a) Top row: Peri-stimulus time histograms (PSTHs) for the "e" key with (top) and without (bottom) RTN for two users. Each PSTH shows the spectral features derived from the left hand, with spectrograms from the 16 electrodes concatenated together. RTN mitigates the significant differences in across-user feature scale and bias. b) Top Row: PSTHs for the "e" key from 3 training users. Note that some of User 4’s features show similar patterns to User 3, while others show similar patterns to User 5. ACM isolates small feature combinations, which are more often shared across users.
Insight 1: sEMG features should be causally normalized per session. sEMG amplitudes routinely differ by an order of magnitude across participants or sessions (3; 8). Figure 1a highlights this disparity: the same channels exhibit both higher baselines and greater variance for User 2 than for User 1. Hence, EMG signals should obviously be normalized between users. However, the method of normalization is critical. Sivakumar et al. (14) apply batch normalization that computes mean and variance over the mini-batch, time, and per-electrode frequency bins. Because each mini-batch mixes data from many users and sessions, these statistics fail to place features from different users/sessions in a common space during training and remain fixed at inference. We instead employ Rolling Time Normalization (RTN), a causal z-scoring of every input feature that updates its mean and variance online. Though simple, this mitigates the scale and shift differences across users, visualized in Figure 1a. Further, RTN requires no calibration and adds virtually no latency for causal, real-time inference. We empirically find this significantly improves zero-shot and finetuned performance.
Insight 2: sEMG feature subsets may be conserved across users. Through visualizing sEMG activity, we empirically find that while two typists rarely share identical high‑dimensional EMG signatures for the same keystroke, they often share subsets of electrodes or frequency bands that behave similarly. A representative example is shown in Figure 1b, where we present the PSTHs of sEMG activity from three users. While the full set of features looks fairly different across users, randomly selected subsets of features have a greater chance of being similar between some users (Augmented View 1 and 2 in Figure 1b). Decoding strategies that hinge on these low‑order combinations therefore stand a better chance of transferring across users. While there are many ways to encourage this, we focus on substantially modifying the hyperparameters of dataset augmentation to achieve this, incurring no test-time computational expense. In particular, we do Aggressive Channel Masking (ACM) that on average zeros out more than half of the electrodes and removes broad spectral chunks from the rest. ACM encourages the model to rely on smaller, more universal feature sets (Figure 1b), which empirically improves zero-shot generalization. We emphasize that this ‘low-order similarity’ perspective is a working hypothesis rather than a definitive statement about the structure of EMG representations: the empirical improvements we observe are consistent with this interpretation but do not prove it.
Insight 3: an inductive bias for bilateral typing. Keyboard typing is fundamentally bilateral: each keystroke is driven solely by the muscles of its own wrist, while the neuromuscular mapping from activation to finger motion is almost mirror‑symmetric across hands. We therefore propose a Split‑and‑Share architecture that accounts for these two facts in the model. Separate, but weight‑shared, subnets encode left‑ and right‑hand sEMG streams in parallel, converging only at the final linear layer where a single vocabulary‑level softmax decodes the character (Figure 2). This enforces hand‑specific locality, but prevents the encoder from entangling spurious cross‑hand correlations. Further, re-use of parameters reduces resource demands: the compact Split‑and‑Share model uses 0.25× the parameters and approximately $60\%$ of the FLOPs of a conventional joint‑hand encoder while matching zero‑shot accuracy.
Summarily, our new, lightweight architecture built on simple modifications from these three insights reduces zero-shot generalization by $31\%$ and finetuning performance by $21\%$ over the prior state-of-the-art in Sivakumar et al. (14). Together, we both increase performance while reducing computational cost, providing a more feasible path towards on-device, performant computation for sEMG interfaces.
Figure 2: a) The bilateral structure of keyboard typing. b) The Split-and-Share macro-architecture.
3 Related Works
Per‑Session Normalization in EMG: The EMG and brain-computer interface (BCI) literature commonly normalize input signals prior to decoding to account for variability. For instance, prior EMG studies expressed amplitudes as a percentage of a reference contraction (maximum voluntary contraction) or applied z-score normalization based on a calibration recording (15). These normalizations enable comparisons across muscles and subjects (3; 8), and several works have normalized EMG features per session for machine learning pipelines (15; 10; 7; 20; 16; 6). In BCI literature, Willett et al. (19) show that continuously updating z‑score statistics on spiking features within each session is essential for their high‑performance speech neuroprosthesis, and Jarosiewicz et al. (5) used online bias correction to improve a point-and-click BCI. As described in Section 2, an important insight is not whether sEMG activity should be normalized, but how it is normalized. We ultimately find RTN significantly outperforms batch normalization.
Masking-Based Regularization for sEMG: SpecAugment, first introduced for speech recognition, has recently been ported to EMG decoding pipelines (12; 14). The version used by Sivakumar et al. (14) is relatively mild: on average only $\sim 6\%$ of inputs are masked per training sample, with a given electrode’s spectrogram never being masked across more than $\sim 24\%$ of the frequencies. ACM corresponds to hyperparameters that are intentionally harsher, with at least half of the electrodes completely blanked in most mini-batches and $\sim 55\%$ of input features masked on average. This encourages the model to infer keystrokes from small, overlapping subsets of channels, discouraging brittle reliance on the full spatiotemporal feature pattern.
Multi-stream architectures and weight sharing: The concept of processing multiple information streams independently before later integration has been established in Automatic Speech Recognition (ASR), where multi-stream systems effectively handle speech at various resolutions or from multiple arrays (9). Similarly, in Sign Language Recognition (SLR), previous works have implemented separate processing streams for each hand (11; 17). Leveraging anatomical symmetry through weight sharing has shown success in pose estimation (21) and SLR (17). In the domain of EMG-based hand-pose recognition, Salter et al. (13) combined data from both hands into a unified dataset for single-model training, though this addresses an inherently unimanual task. For bimanual keystroke decoding, past methods have processed signals from both hands jointly (14; 18; 1). Our Split‑and‑Share architecture keeps the two wrists separate to respect biomechanics, yet ties the weights, effectively training a single-hand encoder with data from both hands. To our knowledge, this has not been explored in EMG keyboard decoding.
4 Methods
4.1 Reduced Spectral Granularity
Let $x_{c}(t)$ denote the raw 2 kHz EMG signal from electrode $c$. Following Sivakumar et al. (14) we first form a log‐power spectrogram
$S_{c}(t,f)=\log_{10}\!\Bigl(\bigl\lvert\operatorname{STFT}(x_{c})\bigr\rvert^{2}_{t,f}+10^{-6}\Bigr),\qquad f=1,\dots,33,$
using a 64‐point FFT ($33$ linearly spaced bins up to 1 kHz) and a hop size of 16. We instead aggregate the 33 bins into six broader, roughly log‐spaced bands, using the same frequency ranges as in (2). We refer to this aggregation as reduced spectral granularity (RSG).
$\mathcal{B}_{1}=[31.25,62.5],\;\mathcal{B}_{2}=[62.5,125],\;\mathcal{B}_{3}=[125,250],\;$
$\mathcal{B}_{4}=[250,375],\;\mathcal{B}_{5}=[375,687.5],\;\mathcal{B}_{6}=[687.5,1000]\;\text{Hz}.$
For each electrode the reduced spectrogram is obtained by
$R_{c}(t,b)=\sum_{f\in\mathcal{B}_{b}}S_{c}(t,f),\qquad b=1,\dots,6.$
Hence, we reduce the spectral dimensionality of each electrode from 33 to 6. This represents an over five-fold reduction in features, and empirically equals or improves performance. This also leads SpecAugment to mask significantly more frequencies, even prior to ACM.
4.2 Rolling Time Normalization
Sivakumar et al. (14) normalizes each electrode’s spectrogram with batch‑norm over the entire mini‑batch, frequency, and time axes. This collapses all features recorded on an electrode into a single distribution and, crucially, re‑uses statistics gathered during training at inference time—an issue when the test session (or user) drifts from the training distribution. RTN replaces this with a causal, per‑feature normalizer that is computed independently for every sample, band, electrode channel and frequency bin. Let
$x_{t,n,b,c,f}\in\mathbb{R},\qquad t=0\dots T-1,\;n=0\dots N-1,$
denote the log‑power value at time‑step $t$, mini‑batch index $n$, band $b$, electrode $c$ and frequency bin $f$. RTN maintains cumulative statistics
$\mu_{t}=\frac{1}{t+1}\sum_{s=0}^{t}x_{s,n,b,c,f},\qquad\sigma_{t}=\sqrt{\left(\frac{1}{t+1}\sum_{s=0}^{t}x_{s,n,b,c,f}^{2}\right)-\mu_{t}^{2}+\varepsilon},$
which are cheap to update via running sums in streaming settings. During a warm‑up period of $T_{w}=125$ frames (the first $1$ second) the statistics are frozen to those computed over the entire warm‑up window. The normalized output is
$\hat{x}_{t,n,b,c,f}=\frac{x_{t,n,b,c,f}-\tilde{\mu}_{t}}{\tilde{\sigma}_{t}},$
where $\tilde{\mu}_{t}$ and $\tilde{\sigma}_{t}$ denote the warm‑up‑frozen statistics for $t<T_{w}$ and the cumulative statistics above for $t\geq T_{w}$. RTN adapts continuously to session-specific non-stationarities, avoids batch-level or training-level statistics, and is causal.
4.3 Aggressive Channel Masking
We apply ACM on the reduced spectral granularity features with $B=6$ bands. We mask inputs in $\nicefrac{{2}}{{3}}$ of mini-batches. For each mini-batch we draw the number of frequency masks, $n_{f}\sim\operatorname{Unif}\{0,1,2\}$. Each mask is then sampled independently for every electrode and sample in the batch:
  1. 1.
    Width $w\sim\operatorname{Unif}\{0,\dots,f_{\max}-1\}$, with $f_{\max}{=}12$. The implementation clamps the width to the number of bands, $w\leftarrow\min(w,B)$. Thus the probability that a single mask erases the entire electrode is $\Pr[w=B]=\tfrac{f_{\max}-B}{f_{\max}}=0.5$.
  2. 2.
    Start index $f_{0}\sim\operatorname{Unif}\{0,\dots,B-w\}$.
  3. 3.
    Set $X_{t,c,f_{0}:f_{0}+w-1}\leftarrow 0$ for all timesteps $t$.
An individual mask already removes a channel with probability $0.5$, while two masks remove a channel with probability $0.75$. Remaining electrodes lose large spectrally-coherent chunks. By forcing the network to succeed even when so many electrodes are fully masked, the augmentation discourages brittle high-order feature dependencies and promotes low-order motifs that transfer across users.
4.4 EMG Encoder Architectures
Following spectrogram normalization, all architectures apply a two-stage encoder composed of a Rotation-Invariant MLP and a Time-Depth Separable Convolution (TDSConv) stack, with a linear character prediction layer following the last TDSConv block. These architectures were first presented in Sivakumar et al. (14) and are reproduced here for completeness.
Rotation Invariant MLP.
Let $\mathbf{x}\in\mathbb{R}^{T\times N\times B\times C\times F}$ denote the input, where $T$ is the number of time steps, $N$ is the batch size, $B=2$ denotes the number of bands (hands), $C=16$ is the number of electrode channels per band, and $F$ is the number of spectral frequency bins. Each band is passed through a Rotation-Invariant MLP, which applies a multi-layer perceptron to rotated versions of the electrode channels to ensure robustness to cyclic spatial shifts:
$\mathbf{h}_{b}=\frac{1}{|\mathcal{O}|}\sum_{o\in\mathcal{O}}\text{MLP}(\text{roll}(\mathbf{x}_{b},o)),\quad\mathcal{O}=\{-1,0,1\},$
where each rotated input is flattened across $C\times F$ before the MLP is applied. This results in an embedding $\mathbf{h}_{b}\in\mathbb{R}^{T\times N\times D}$ per band. Concatenating the left and right hand features yields a total input of shape $\mathbf{h}\in\mathbb{R}^{T\times N\times 2D}$ to the TDSConv blocks in the baseline model, whereas each hand is processed independently in the Split-only and Split-and-Share variants.
TDSConv block architecture.
The TDSConv stack (4) is composed of alternating temporal convolution and fully connected blocks. Each convolutional block first reshapes the input from $\mathbb{R}^{T\times N\times D}$ to $\mathbb{R}^{N\times K\times H\times T}$, where $D=K\cdot H$ denotes the total feature dimension, $K$ is the number of convolution channels, and $H$ is the per‑channel hidden width. A 2‑D convolution with kernel size $1\times w$ is then applied along the time axis:
$\mathbf{z}[n,k,h,t]=\sum_{i=0}^{w-1}\sum_{k^{\prime}=1}^{K}\theta_{k,k^{\prime},i}\;\mathbf{h}[n,k^{\prime},h,t-i],\qquad\theta\in\mathbb{R}^{K\times K\times 1\times w}.$
Because the kernel height is 1, the same weights are reused at every hidden‑width position $h$. The convolution still mixes all $K$ input channels to produce each of the $K$ output channels, giving the layer $K^{2}w$ parameters. The output is passed through ReLU, summed with the residual, and normalized using LayerNorm. The subsequent fully connected block consists of two linear layers with a ReLU in between and a residual connection:
$\mathbf{z}_{\text{fc}}=\text{LayerNorm}(\text{FC}_{2}(\text{ReLU}(\text{FC}_{1}(\mathbf{z}_{\text{conv}})))+\mathbf{z}_{\text{conv}}).$
Architecure variants.
In the architecture of Sivakumar et al. (14), which we refer to as the Joint-Hand architecture, both hands are processed jointly following the Rotation-Invariant MLP, yielding a post-MLP embedding of $2D=768$. The TDSConv stack operates with this full dimensionality throughout. In contrast, both the Split-only and Split-and-Share models operate on embeddings of $D=384$ per hand, and apply the TDSConv blocks separately to each hand’s input. This reduces the width of all fully connected layers from $768\times 768$ to $384\times 384$, leading to a four-fold reduction in parameters per FC layer, while preserving expressiveness by keeping the number of convolution channels $K=24$ and kernel width $w=32$ unchanged. Despite duplicating the encoder, the Split-only model has lower total FLOPs—approximately 60% of the baseline—due to the narrower per-hand embeddings. The Split-and-Share model uses the same dual-stream structure but shares all encoder weights between hands. This model has roughly half the parameters of the Split-only model, since almost all parameters (excluding the final prediction layer) are shared between the two encoders. Notably, the Split-only and Split-and-Share models have the same FLOPs.
Figure 3: EMG encoder architectures. Left (blue) and right (red) hand specific modules only process inputs from a single hand, with hand-specific weights. Joint (purple) modules jointly process inputs from both hands. Shared (gray) modules process inputs from either hand using identical weights. a) Joint-Hand baseline architecture of Sivakumar et al. (14), b) Split-only architecture, in which hand-specific modules process signals from each hand separately. c) Split-and-Share architecture, where shared-weight modules process signals from each hand separately.
To evaluate whether representational capacity might further enhance the strong performance of the Split-and-Share architecture, we also test an Upscaled Split-and-Share variant. This model increases the per-hand embedding size to $D=528$ and expands the number of convolution channels in the final two TDS blocks from 24 to 48. Nonetheless, the total number of parameters remains about half that of the baseline, and the FLOPs are only modestly higher (about 15%). We refer to the Upscaled Split-and-Share model as SplashNet and the smaller Split-and-Share model as SplashNet-mini
4.5 Backspace-aware beam search with 6-gram Character LM
For decoding we adopt the backspace‑aware beam search of Sivakumar et al. (14), which keeps the 50 most‑probable CTC label prefixes at each frame. Blank transitions simply update the score of an unchanged prefix, while the highest‑scoring non‑blank characters extend each hypothesis and are re‑ranked with a 6-gram character LM prior. When a backspace symbol appears, the algorithm retracts the last language‑model contribution, letting deletions correct earlier mistakes without extra passes. After the final frame any open LM context is closed, and the best‑scoring prefix provides the decoded keystroke sequence. All beam-search parameters are the same as in Sivakumar et al. (14)
4.6 Dataset Splits
The official emg2qwerty protocol introduced by Sivakumar et al. (14) assigns 100 participants to the training pool and eight to a held-out test pool. For the 96 training participants with $\geq 4$ recording sessions, every session except the final two is used to train generic models, while those last two sessions form the authors’ validation set; models are selected on this set before being evaluated on the eight held-out participants.
Because the validation data come from the same individuals seen during training, this procedure encourages models that memorize participant-specific idiosyncrasies rather than ones that generalize across users. To illustrate this effect while limiting computational cost, we evaluate training domain validation performance using the validation sessions from 18 of the 96 training users, which we term the training domain validation set. The same 18 users are also used in Appendix A.8, where they are fully held out from training to provide a more appropriate validation signal.
To obtain a validation signal that better predicts zero-shot performance in the main experiments, we instead exploit the four training-pool participants who recorded <4<4 sessions and were therefore excluded from model fitting in the original setup. We take the final session of each of these four participants and combine them into what we term the other domain validation set. These participants exhibit noisier EMG data than those in the official test pool, providing a harder proxy for real-world generalization. Nonetheless, they provide a way of validating zero-shot generalization without requiring a complete restructuring of the official dataset splits.
5 Results
5.1 Zero-shot model performance
Table 1: Zero-shot CER (%, mean $\pm$ s.d. across participants), GFLOPs, and parameters. Columns in gray correspond to training domain validation results, which are reported for transparency but not used as indicators of generalization.
Method Train domain val
Other
domain
val
Test
domain
val
Test
domain
test
GFLOPs
(30 s)
Params
Sivakumar et al. 2024 $12.14$ $72.07$ $52.10\pm 5.54$ $51.78\pm 4.61$ 61.61 5.29M
+ RSG $13.52$ $67.48$ $47.26\pm 5.26$ $47.18\pm 5.19$ 54.15 4.96M
+ RTN $13.09$ $61.95$ $39.49\pm 7.45$ $39.15\pm 6.20$ 54.15 4.96M
+ ACM $23.47$ $63.08$ $42.62\pm 7.18$ $42.62\pm 7.10$ 54.15 4.96M
+ RTN + ACM $21.71$ $58.85$ $36.41\pm 7.21$ $36.42\pm 7.11$ 54.15 4.96M
+ Split $23.93$ $58.64$ $28\pm 6.91 37.37$$37\pm 7.34$ 36.84 2.68M
   + Share (SplashNet-mini) $26.44$ $58.20$ $36.46\pm 7.09$ $36.41\pm 7.30$ 36.84 1.38M
   + Upscale (SplashNet) $20.59$ $\bf{56.95}$ $\bf{35.49\pm 7.56}$ $\bf{35.67\pm 6.79}$ 71.38 2.58M
Before turning to the main cross-user generalization results, we first highlight a key discrepancy between training domain validation and held-out user evaluation. As shown in Table 1, the CERs on the training domain validation set are not only lower, but the relative ranking of models follows a markedly different pattern than on held-out user evaluation sets. This illustrates why training domain validation provides a misleading signal for cross-user generalization and should not be used for model selection. In Appendix A.8, we further show that when these same 18 users are held out entirely, their CERs rise substantially and align closely with the held-out evaluation trends shown here.
Our analysis centers on beam-search decoding using the same 6-gram character LM as in Sivakumar et al. (14); corresponding greedy-decoding results appear in Appendix A.3. All reported p-values were obtained from one-tailed paired t-tests across participants.
First, we apply RSG to replace the 33-bin spectrogram with six coarser frequency bands. RSG leads to a modest but consistent improvement in CER from $51.78\%$ to $47.18\%$ (p=7ep=7e-55). Part of this performance improvement may stem from the implicit amplification of frequency masking under SpecAugment (due to the lower spectral resolution).
Second, we replace batch-level normalization with RTN. RTN significantly improves zero-shot generalization (p=6ep=6e-44), reducing the CER to $39.15\%$. This suggests that input scale and shift differences across users are a primary obstacle to cross-user generalization in EMG decoding.
Third, we apply ACM, which encourages the model to rely on lower-order combinations of input features. Without RTN, ACM reduces zero-shot CER to $42.6\%$ ($p=.02$ vs. +RSG only). Combined with RTN, ACM reduces zero-shot CER to $36.42\%$ (p=5ep=5e-33 vs. +RSG+RTN), providing a strong Joint-Hand baseline.
Fourth, we explore architectural modifications that better reflect the causal and bilateral structure of EMG typing. We evaluate a Split-only model that encodes each hand separately without parameter sharing, achieving a CER of $37.37\%$. Despite having only half the parameters and $66\%$ of the FLOPs of our Joint-Hand baseline, this model performs competitively, though its lack of shared parameters may limit data efficiency. We then evaluate SplashNet-mini, where both hands are encoded via identical weight-shared encoders. This model achieves a $36.41\%$ CER—on par with the Joint-Hand baseline—while using just a quarter of its parameters and $66\%$ of the FLOPs. Finally, by increasing the embedding width and expanding the final convolutional layers, we create SplashNet, with similar FLOPs to the baseline of Sivakumar et al. (14) but still half the parameters. This yields a further improvement, reducing the CER to $35.67\%$ ($p=.049$ vs. SplashNet-mini).
Together, these results demonstrate that a combination of architectural priors, per-session normalization and principled regularization can significantly improve zero-shot EMG decoding.
5.2 Finetuned Model Performance
Table 2: Finetuned CER (%, mean $\pm$ s.d. across participants), GFLOPs, and parameters.
Method
Test
domain
val
Test
domain
test
GFLOPs
(30 s)
Params
Sivakumar et al. 2024 8.31 $\pm$ 3.19 6.95 $\pm$ 3.61 61.61 5.29M
+ RSG 6.70 $\pm$ 3.22 6.92 $\pm$ 3.79 54.15 4.96M
+ RTN 6.47 $\pm$ 2.92 6.63 $\pm$ 3.58 54.15 4.96M
+ ACM 7.18 $\pm$ 3.33 7.45 $\pm$ 3.87 54.15 4.96M
+ RTN + ACM 6.55 $\pm$ 2.91 6.53 $\pm$ 3.27 54.15 4.96M
+ Split 6.04 $\pm$ 2.81 6.10 $\pm$ 3.32 36.84 2.68M
    + Share (SplashNet-mini) 6.13 $\pm$ 2.96 5.87 $\pm$ 3.04 36.84 1.38M
    + FT Unshare 5.85 $\pm$ 2.83 5.96 $\pm$ 3.28 36.84 2.68M
    + Upscale (SplashNet) 5.46 $\pm$ 2.60 5.51 $\pm$ 2.81 71.38 2.58M
     + FT Unshare 5.57 $\pm$ 2.65 5.67 $\pm$ 2.97 71.38 5.06M
We next evaluate the performance of models finetuned on user-specific data. As in Sivakumar et al. (14), we maintain identical training hyperparameters during both generic pretraining and finetuning. While this simplifies the analysis, we note that the optimal hyperparameters for finetuning likely differ from those used during pretraining, and further gains may be achievable through phase-specific hyperparameter tuning. We also note that, because each recording session involved doffing and re-donning the wristbands, the finetuning experiments inherently probe generalization across electrode placements in different sessions.
We report all results using beam search with an external character-level language model (LM), which remains the standard for achieving state-of-the-art performance in CTC-based decoding pipelines. While our models consistently outperform the baseline of Sivakumar et al. (14) under beam search, we observe slightly worse performance under greedy decoding. We attribute this in part to the more aggressive channel masking induced by our reduced spectral representation (§4), which we discuss further in Appendix A.3.
First, we assess whether the same methods that improved zero-shot generalization also enhance performance in the finetuned setting. Using RSG alone yields similar performance to the baseline of Sivakumar et al. (14), with a CER of 6.91%, while adding RTN yields a significantly improved CER of 6.63% ($p=.028$ vs. +RSG). Interestingly, applying ACM in isolation—without RTN—leads to worse performance than the baseline of Sivakumar et al. (14), whereas the Joint-Hand baseline model, with both ACM and RTN, matches the model with RTN alone (6.53% CER). This suggests that masking-induced variability may hinder learning when the model lacks an appropriate normalization strategy to stabilize the input feature space. Moreover, unlike the generic case, finetuned models may rely more on higher-order feature correlations that are relatively stable across sessions for the same user, diminishing the benefits of aggressive masking.
We next examine architectural changes that explicitly encode inductive biases about the bilateral and causal structure of EMG typing. Both the Split-only and SplashNet-mini models yield substantial improvements over the Joint-Hand baseline, with CERs of 6.10% and 5.87% (p=6ep=6e-33 and p=2ep=2e-33, respectively). For the latter, we evaluate two strategies during finetuning: either maintaining shared weights or duplicating them to allow separate adaptation per hand. Interestingly, both strategies yield similar performance, suggesting that hand-specific encoder weights are unnecessary even in the finetuned setting, although it is possible that unsharing weights could become advantageous with more user-specfic data for finetuning.
Finally, we evaluate the SplashNet model. Again, we do not find any benefit from unsharing the weights during finetuning. With the weights kept shared, SplashNet achieves the best performance overall with a CER of 5.51% ($p=.02$ vs. SplashNet-mini), establishing a new state-of-the-art for user-specific EMG keystroke decoding in this benchmark.
Figure 4: Zero-shot and finetuned CER distribution across users. Each of the 8 test users are represented by a dot, with lines connecting the same user across models. Boxplots depict median and interquartile ranges. Our methods improve performance for all participants relative to the baseline of Sivakumar et al. (14), with some participants showing very large improvements: two users reach CER between 20-30% in the zero-shot setting, and one user attains a CER below 2% when finetuned.
6 Discussion
The central contribution of this work is instilling simple, well-motivated priors through preprocessing, augmentation and architecture that can close a surprising fraction of the generalization gap in wrist‑EMG typing. SplashNet‑mini and SplashNet both achieve large absolute and relative CER reductions (15.4‑15.4 pp / $31.1\%$ zero‑shot; 1.44‑1.44 pp / $20.7\%$ after fine‑tuning) while cutting parameters to ¼–½ of the baseline and FLOPs to 0.6–1.15×. These gains are on par with (and complementary to) the $\sim 25\%$ error reduction CTRL-Labs at Reality Labs et al. (2) reported from doubling dataset size in a handwriting task, suggesting that principled inductive biases are as potent as raw data scaling for sEMG.
A practical ambition articulated in Sivakumar et al. (14) is to run the entire decoder on the wristbands themselves, thereby mitigating concerns around latency, privacy, and robustness to Bluetooth interference. Achieving on‑band inference will require an architecture that is not only light‑weight but also fully split—able to process and output keystrokes from each wrist independently, without any cross‑hand coupling. SplashNet moves part‑way toward this goal by duplicating (and sharing) the encoder streams, yet it still merges information at the final linear output layer, so each wrist must communicate its embeddings to the other. Removing this last dependency, or replacing it with a low‑bandwidth handshake, remains an open engineering challenge and a fruitful direction for future model architecture work. Another promising direction is exploring hybrid cross-hand architectures that retain split-and-share components but introduce limited, structured interactions between the hands, potentially offering a middle ground between the Split-and-Share and Joint-Hand extremes.
Beyond these concrete results, our study points to several avenues. Normalization: RTN is a single‑pass causal z‑score; richer adaptive schemes (momentum updates, learnable affine transforms, or brief self‑supervised calibration) may yield further robustness. Structured masking: ACM already improves zero‑shot CER, but spatially contiguous electrode “drop‑blocks”, global time masks, or anatomical adjacency priors could guide the network toward even more transferable features. Model scale and self‑supervision: SplashNet shows that capacity can be reinvested profitably once good priors are in place; coupling larger split‑and‑share encoders with masked‑prediction pre‑training may unlock still‑higher accuracy without additional labels.
Two limitations deserve emphasis. First, our models are developed for the practical beam-search + LM pipeline; under pure greedy decoding, the Joint-Hand baseline scores the best on unseen users, and ACM dents finetuned accuracy (details and discussion in Appendix A.3). This likely reflects the fact that the Joint-Hand baseline can exploit cross-hand co-articulatory patterns, which tend to correlate strongly with character bi- and tri-grams. These weak statistical regularities effectively act as an implicit language model under greedy decoding. By contrast, Split-and-Share models do not have access to these bilateral dependencies, and ACM further suppresses within-user cues by enforcing lower-order feature reliance. Once an external 6-gram LM is applied, however, its much stronger prior readily compensates for both the lost co-articulatory signal of the Split-and-Share models and the reduced within-user discriminability induced by ACM, thereby revealing the underlying generalization advantage of our approach. Second, all experiments use healthy participants; the extent to which our priors transfer to populations with motor or limb differences (where electrode placement and muscle recruitment differ markedly) remains to be tested.
In sum, we show that the long‑standing vision of “keyboard‑quality” EMG typing can be advanced not only by more data but also by better assumptions. By pairing causal normalization, aggressive yet structured regularization, and a symmetry‑aware encoder, we make significant strides towards truly out‑of‑the‑box wrist‑sEMG typing—paving the way for real‑world assistive and AR/VR interfaces.
7 Acknowledgments
This work was supported by the following awards to JCK: NSF CAREER 1943467, NIH DP2NS122037, NIH R01NS121097. The authors would also like to thank Andrea Ortone for carefully checking the FLOPs calculations and providing helpful feedback.
Appendix A Technical Appendices and Supplementary Material
A.1 Training Details
All models were trained for 150 epochs with the hyperparameter configuration of Sivakumar et al. [2024]. we used the same optimizer (Adam), the same LR schedule (linear warmup from 1e-8 to 1e-3 for first 10 epochs, followed with cosine annealing to 1e-6 until epoch 150), and the same augmentations. These included a rotation augmentation, which shifts all electrodes on each band one electrode to the left or right for each training sample, and a temporal alignment jitter augmentation, which jitters the EMG signals from each hand by a maximum offset of 60 ms. We only made two modifications to hyperparameters:
  1. 1.
    Input window length. We employ 16 s training windows, rather than the 4 s clips used by Sivakumar et al. [2024]. This was done to speed up training.
  2. 2.
    SpecAugment settings.
    • With ACM. We raise the maximum frequency mask width to 12 and disable time masking.
    • Without ACM. We retain the original SpecAugment parameters of Sivakumar et al. [2024] (maximum frequency mask width = 4; up to 3 time masks per electrode per sample, each masking out up to 200 ms).
A.2 Evaluation Details
All validation and test sessions were loaded as a single sample (i.e. with batch size of 1). By comparison, Sivakumar et al. [2024] loaded validation (but not test) sessions as shortened chunks. Although this is unlikely to have affected their greedy decoding results as their architecture (and ours) has a receptive field of only 1 second, this might have affected the validation performance they reported with an external LM, which would lose the character history between each chunk and therefore make poorer predictions. This might explain why the validation CER they report with beam-search (8.31% CER) is conspicuously higher than their test performance (6.95% CER) compared to the gaps we see between validation and test sessions.
A.3 Greedy vs. Beam‑Search Decoding
Our main text focuses on beam‑search decoding with an external 6‑gram language model (LM), because this is the configuration most relevant to realistic, latency‑constrained deployment. For completeness, we also report greedy CTC decoding results,i.e. decoding without any LM.
Table 3: Zero-shot CER (%, mean across participants). Columns in gray correspond to training domain validation results, which are reported for transparency but not used as indicators of generalization. Note that the training domain validation results shown here with greedy decoding were computed over all 96 training users; the beam search decoding results are computed on the 18 user subset.
Method Train domain val (greedy) Train domain val (beam)
Other
domain
val
Other
domain
val
Test
domain
val
Test
domain
val
Test
domain
test
Test
domain
test
Sivakumar et al. 2024 22.51 12.14 72.44 72.07 55.57 52.10 55.38 51.78
+ RSG 23.59 13.52 68.55 67.48 52.24 47.26 52.27 47.18
+ RTN 23.49 13.09 64.19 61.95 46.31 39.49 46.07 39.15
+ ACM 34.77 23.47 65.60 63.08 48.87 42.62 49.23 42.62
+ RTN + ACM 32.85 21.71 61.82 58.85 43.72 36.41 43.80 36.42
+ Split 35.59 23.93 61.74 58.64 45.73 37.28 45.69 37.37
+ Share (SplashNet-mini) 37.72 26.44 61.07 58.20 45.33 36.46 45.26 36.41
   + Upscale (SplashNet) 33.57 20.59 60.16 56.95 44.79 35.49 44.78 35.67
Zero‑shot generalization.
Table 3 shows that, in the zero‑shot regime (i.e. on non-training domains), the joint‑hand architecture with RTN and ACM often surpasses the Split‑and‑Share architecture under greedy decoding, even though the ranking reverses once beam search is applied. The most plausible explanation is that the joint‑hand encoder can observe both wrists simultaneously and therefore learns a stronger implicit LM, capturing bi‑ and tri‑gram dependencies spread across hands. While that emergent linguistic prior is helpful when no external LM is available, it becomes redundant—and potentially counter‑productive—once a more reliable 6‑gram LM is introduced at inference time.
Table 4: Finetuned CER (%) with and without beam search.
Method
Test
domain
val
Test
domain
val
Test
domain
test
Test
domain
test
Sivakumar et al. 2024 11.39 8.31 11.28 6.95
+ RSG 12.11 6.70 12.50 6.92
   + RTN 11.83 6.47 11.75 6.63
   + ACM 14.24 7.18 14.74 7.45
   + RTN + ACM 12.70 6.55 12.90 6.53
    + Split 12.67 6.04 12.89 6.10
    + Share 13.01 6.13 13.22 5.87
     + FT Unshare 12.68 5.85 13.07 5.96
     + Upscale (SplashNet) 12.10 5.47 12.39 5.51
     + FT Unshare 11.80 5.57 12.13 5.67
Fine‑tuned models.
A different pattern emerges after user‑specific fine‑tuning (Table 4). None of our variants fully match the baseline of Sivakumar et al. [2024] under greedy decoding, despite decisive gains once beam search is enabled. Because the only changes between the baseline and the "+ RSG" model are (i) four‑fold longer clips during training, (ii) coarser spectral resolution, and (iii) a five‑fold higher probability of channel masking, the degradation must stem from one (or a combination) of these factors.
Two clues implicate aggressive channel masking (ACM). First, the model with RTN but not ACM achieves the best greedy scores, whereas the model with just ACM but not RTN yields the worst greedy performance. Second, the performance gap between the model with just RTN and the model with both RTN and ACM vanishes under beam search, indicating that the external LM compensates for information lost when ACM forces the network to rely on low‑order feature combinations. We therefore hypothesize that ACM, while beneficial for cross‑user generalization, removes within‑user cues that help distinguish confusable keystrokes, a weakness that beam‑search decoding can largely recover.
To confirm that the implicitly increased channel masking in the "+ RSG" model (and all other models) is largely responsible for the uniformly worse greedy decoding performance we see compared to the baseline of Sivakumar et al. [2024], we ran an additional experiment in which we trained a model similar to our +RSG model, but apply SpecAugment on the 33-bin spectrograms from each channel before the bin aggregation of RSG (rather than after). This keeps the extent of masking equivalent to that of Sivakumar et al. 2024 while also allowing us to use the RSG frontend. This model does not perform significantly worse than the baseline of Sivakumar et al. 2024 with greedy decoding, confirming our suspicion that the core reason for the worse greedy decoding results in the finetuned case is increased channel masking.
Table 5: Finetuned CER (%) under greedy decoding. Values are mean $\pm$ standard deviation.
Method
Test
domain
val
Test
domain
test
Sivakumar 2024 $1.39\pm 4.28 11$$1.28\pm 4.45$
+ RSG $12.11\pm 4.67$ $12.50\pm 4.97$
+ RSG w/ pre-aggregation masking $1.02\pm 4.32 11$$1.53\pm 4.70$
In summary, greedy decoding accentuates two complementary inductive biases: (1) joint‑hand encoders learn a useful internal LM, an advantage that vanishes once an external LM is applied, and (2) aggressive masking of input channels trades cross‑user generalization for within‑user discriminability, a trade‑off that RTN and beam‑search decoding can effectively offset.
A.4 Calculation of FLOPs
FLOPs were measured using FlopTensorDispatchMode in PyTorch with an arbitrary 30-second input.
A.5 Additional Ablations on ACM Masking, RTN Sliding Windows, and RSG
Table 6: Zero-shot CER (%, mean $\pm$ s.d. across participants) for ablations on ACM mask width and RTN sliding window. “-RSG” corresponds to applying ACM on the full-resolution spectrogram as described in the text.
Method
Other
domain
val
Test
domain
val
Test
domain
test
SplashNet-Mini $58.20\pm 10.50$ $36.46\pm 7.09$ $36.41\pm 7.30$
+ 4s SW inference $60.43\pm 7.86$ $37.57\pm 7.18$ $37.71\pm 7.47$
+ 4s SW train + inference $60.43\pm 7.70$ $36.95\pm 7.89$ $36.69\pm 7.80$
+ 16s SW inference $57.24\pm 11.00$ $36.43\pm 7.30$ $36.73\pm 7.40$
+ mask width = 8 $58.12\pm 11.72$ $36.57\pm 7.02$ $36.05\pm 6.74$
+ mask width = 16 $57.65\pm 10.47$ $36.66\pm 7.21$ $36.92\pm 7.40$
- RSG $60.70\pm 11.23$ $37.90\pm 6.58$ $37.19\pm 6.33$
Joint-Hand baseline (+ RSG + RTN + ACM) $58.85\pm 10.50$ $36.41\pm 7.21$ $36.42\pm 7.11$
+ mask width = 8 $59.41\pm 10.63$ $37.55\pm 6.97$ $37.30\pm 6.90$
+ mask width = 16 $57.17\pm 11.32$ $36.95\pm 7.62$ $36.47\pm 7.96$
Table 7: Finetuned CER (%, mean $\pm$ s.d. across participants) for ablations on ACM mask width and RTN sliding window.
Method
Test
domain
val
Test
domain
test
SplashNet-Mini $6.13\pm 2.96$ $5.87\pm 3.04$
+ 4s SW inference $6.10\pm 2.91$ $6.28\pm 3.09$
+ 4s SW train + inference $6.34\pm 2.86$ $6.37\pm 3.10$
+ 16s SW inference $5.76\pm 2.64$ $5.74\pm 2.86$
+ mask width = 8 $5.85\pm 2.72$ $6.02\pm 3.11$
+ mask width = 16 $6.09\pm 2.88$ $5.72\pm 2.96$
- RSG $5.87\pm 2.73$ $5.75\pm 3.06$
Joint-Hand baseline (+ RSG + RTN + ACM) $6.55\pm 2.91$ $6.53\pm 3.27$
+ mask width = 8 $6.14\pm 2.85$ $6.56\pm 3.60$
+ mask width = 16 $6.55\pm 3.09$ $6.45\pm 3.18$
We conducted additional ablations to examine the sensitivity of model performance to the strength of ACM, the presence of RSG, and the temporal window used for RTN normalization. Unless otherwise noted, none of the differences reported here reached significance under a two-tailed paired $t$-test across participants.
ACM mask width.
In our main experiments, ACM uses a maximum frequency mask width of 12 bins, while models without ACM use a maximum width of 4. Here, we trained SplashNet-mini and the Joint-Hand baseline with maximum mask widths of 8 and 16. As shown in Tables 6 and 7, these variations produced small, nonsignificant changes in CER in both the zero-shot and finetuned settings, indicating that performance is not particularly sensitive to the precise mask width within this range.
-RSG.
We also evaluated a variant without RSG, in which ACM was applied to the full-resolution spectrogram by masking a 6-bin dummy vector and inverting the 33-to-6 bin mapping used in RSG to obtain a 33-bin mask. Although performance differences relative to the standard RSG configuration were small and nonsignificant, this variant considerably increases both compute and memory requirements since it has more than sixfold greater input feature dimensionality.
RTN sliding window.
Our original RTN normalization uses all past time points within a sample to compute normalization statistics. We additionally evaluated inference-time sliding-window variants with 4-second and 16-second windows. Using a 4-second window led to a small but statistically significant degradation in zero-shot performance ($p<0.05$, two-tailed), while a 16-second window yielded performance comparable to the default setting. This likely reflects the fact that models were trained on 16-second samples and benefited from matching normalization context at test time. When models were also trained with a 4-second RTN sliding window, this degradation disappeared, and differences were no longer significant in either the zero-shot or finetuned settings.
Summary.
Across all tested configurations, performance differences were small and generally nonsignificant. ACM is robust to maximum frequency mask width changes in the range of 8–16 bins. Removing RSG does not significantly degrade performance but incurs higher compute and memory costs. RTN performance remains stable across temporal windows when training and inference are matched, with only a modest increase in CER when using 4 second windows at inference time alone.
A.6 Aggressive channel masking with mean imputation
Table 8: CER (%, mean $\pm$ s.d. across participants) for different ACM masking configurations.
Method
Test
domain
val
Test
domain
test
Sivakumar et al. 2024 $52.10\pm 5.54$ $51.78\pm 4.61$
+ RSG $47.26\pm 5.26$ $47.18\pm 5.19$
+ ACM $42.62\pm 7.18$ $42.62\pm 7.10$
+ ACM (mean impute) $47.59\pm 7.16$ $48.43\pm 6.24$
+ RTN + ACM $36.41\pm 7.21$ $36.42\pm 7.11$
We performed an ablation to test whether RTN might interact with ACM by stabilizing input features such that the default masking value of 0 corresponds to their per-session mean. When using ACM with RTN, the default mask value of 0 corresponds to the per-sample mean for each feature, whereas under BatchNorm this same value can be far out of distribution for a given feature from a given sample. To test whether it is important that the masking value is the per-sample feature mean, we replaced RTN with standard BatchNorm while setting the ACM masking value to the per-sample mean of each feature, training a Joint-Hand model with RSG, BatchNorm, and ACM using per-channel sample mean imputation.
Somewhat surprisingly, this variant performed on par with the model with RSG alone and substantially worse than the RSG + BatchNorm + ACM configuration with standard zero imputation. This suggests that simply matching the mask value to the per-session mean is not sufficient to reproduce the performance benefits obtained with RSG + RTN + ACM, let alone to obtain the benefits of ACM in the absence of RTN. More generally, these results indicate that the choice of masking value being stable across samples (rather than corresponding to the per-sample mean) appears to be the more important factor for ACM’s effectiveness in this setting.
A.7 UMAP analyses on early intermediate representations
Figure 5: UMAP visualization of model activations after the first TDSConv block for four models (+RSG, +RSG+ACM, +RSG+ACM+RTN, and SplashNet-mini). We extracted activations from every 100th timestep from one session of each of the 8 held-out users. Colors indicate the user identity of each point. In the models without RTN, some users’ representations occupy largely disjoint regions of the activation manifold, whereas models with RTN (+RSG+ACM+RTN and SplashNet-mini) produce markedly more overlapping per-user representations, indicating improved cross-user alignment in the learned feature space.
A.8 Alternative train-test split analyses
Table 9: CER (%, mean $\pm$ s.d. across participants) with a new 78/18 train/validation split. The first column shows performance on the same 18 users when they were included in the original training set (i.e., the previous training domain validation performance), while the remaining columns show results when these 18 users are held out for validation. The test set remains the same 8 held-out users.
Method
Train domain
val (prev. split)
Train domain
val (18 held-out)
Test domain
val (18 held-out)
Test domain
test (18 held-out)
+ RSG $13.52\pm 7.59$ $58.42\pm 10.43$ $47.89\pm 6.36$ $49.01\pm 5.35$
+ RTN $13.09\pm 6.32$ $53.99\pm 11.14$ $42.81\pm 6.75$ $42.62\pm 6.89$
+ ACM $23.47\pm 9.74$ $55.95\pm 10.46$ $44.59\pm 6.47$ $45.10\pm 7.26$
+ RTN + ACM $21.71\pm 9.67$ $51.25\pm 11.49$ $38.72\pm 6.63$ $39.05\pm 7.22$
+ Split $23.93\pm 10.99$ $51.57\pm 11.78$ $39.81\pm 6.72$ $40.14\pm 7.16$
    + Share $26.44\pm 10.64$ $49.84\pm 12.57$ $38.60\pm 7.52$ $38.45\pm 7.86$
A.9 Analyses on emg2pose dataset
Table 10: Mean angular error (degrees) and landmark distance ($m$) under different generalization regimes. All results are in the "tracking" setting.
Generalization Model
Angular
Error
Landmark
Dist
User RTN+ACM 7.5951 10.1567
Stage RTN+ACM 11.1243 15.2404
User, Stage RTN+ACM 10.8433 15.3644
User RTN 7.6007 10.1641
Stage RTN 11.1627 15.2241
User, Stage RTN 10.9085 15.3851
User ACM 7.6388 10.2394
Stage ACM 11.2601 15.4489
User, Stage ACM 11.0343 15.6413
User Baseline 7.6549 10.2585
Stage Baseline 11.2892 15.4360
User, Stage Baseline 11.1222 15.7388
Although the focus of this work is on EMG keystroke decoding, we also performed preliminary experiments to assess how RTN and ACM transfer to the emg2pose benchmark [Salter et al., 2024]. A key consideration is that the baseline model of Salter et al. [2024] uses a learned convolutional featurizer that immediately mixes signals across all electrodes, rather than spectrogram-based per-electrode features as in our keystroke decoding setting. This architectural difference makes a direct application of RTN and ACM less straightforward.
To adapt these methods, we applied RTN and ACM to the outputs of the first convolutional layer in the featurizer. Since these intermediate features are not spectrograms, ACM was implemented by randomly zeroing out 50% of the features in one-third of training samples and 75% in another third. The resulting performance, summarized in Table 10, shows minimal differences relative to the baseline in both mean angular error and landmark distance across user, stage, and combined generalization regimes.
These findings are somewhat unsurprising: applying ACM after early feature mixing likely limits its ability to enforce robust low-order structure, and RTN is less meaningful without clear per-electrode feature boundaries. Future work should explore applying these methods directly to the raw EMG signal before featurization, or using architectures with explicit per-electrode feature streams (e.g., spectrogram-based or structured learned featurizers), where RTN and ACM may have stronger effects.
A.10 Compute Resources
All experiments were run on a single RTX 4090 (with 24GB VRAM) GPU. Each generic model took roughly 24 hours to train. Finetuning took 20-30 minutes for each of the 8 test users, and an additional 30 minutes to evaluate with beam search for each of the 8 test users.
A.11 Dataset License
The emg2qwerty dataset [Sivakumar et al., 2024] is available under the CC-BY-NC-4.0 license, which this work is compliant with.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction claim that Rolling Time Normalization, Aggressive Channel Masking, and Split-and-Share Encoders all help for EMG keystroke decoding, which we show in our results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss limitations in the Discussion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We have no theoretical results. Everything is empirical.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We describe our methods in detail, including exactly how our methods differ from Sivakumar et al. [2024]. We detail our architecture, preprocessing, and augmentation methods, as well as how we developed our models and our dataset splits.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All code is provided in the supplementary material, along with instructions on how to use the code to reproduce our results and evaluate our models.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide all of these details in the Appendix, and state that they were almost all taken from Sivakumar et al. [2024] except for some exceptions, which we provide reasons for.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provide standard deviation values in our tables to give an indication of the distribution of the performance of our models across users, following Sivakumar et al. [2024], and we state explicitly that the metrics in the tables are mean and standard deviation. We also show the per-participant CERs for our key models. Statistical significance is reported using appropriate tests (1-tailed paired t-tests across participants).
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe our compute resources and the time taken for each experiment in the appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The models we propose have no straightforward route to being misused, we do not conduct research with human participants, and we mention the limitations of the dataset we use in that it is not representative of users with motor impairments.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss positive societal impacts, such as always-available typing interfaces for AR/VR. We also make note of concerns around user privacy with such a device that streams biological signals to a computer, which motivates a future direction for our work towards fully on-device keystroke recognition.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We propose methods to improve decoding of keystrokes from EMG wristbands, which we do not believe has any straightforward risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators of the emg2qwerty dataset are explicitly credited and cited throughout the paper. The CC-BY-NC-4.0 license is mentioned in the appendix, and we do not violate its terms.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the code for our models, along with documentation of how to use it.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing or research with human subjects
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects. We use a previously released dataset in compliance with its terms.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs were not used in any of our methods.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.